1) More H2 gas is added to the system at time t2 while the temperature is held constant. Which of the following will most likely occur?

A)The value of the equilibrium constant will decrease

B) The amount of N2 will increase

C) The amount of NH3 will increase

D)The value of the equilibrium constant will increase


2) X(g) + Y(g) ⇄ 2Z(g)

When 6.00 mol each of X(g) and Y(g) are placed in a 1.00 L vessel and allowed to react at constant temperature according to the equation above, 8.00 mol of Z(g) is produced. What is the value of the equilibrium constant, Kc ?

A)36

B)16

C)2

D)8




1
Expert's answer
2021-04-07T11:29:38-0400

1) The value of the equilibrium constant will decrease because the equilibrium constant is expressed in terms of pressure of H2, and the pressure of H2 is in the denominator.

2) 8 mol of Z(g) are in the vessel and 2 mol of both X(g) and Y(g) are left, so, calculate the constant:


"K_c=\frac{[Z]^2}{[X][Y]}=\frac{8^2}{2·2}=16."